The area of a rectangular wall of a barn is 198 square feet. Its length is 4 feet longer than twice its width. Find the length and width of the wall of the barn.
width= ? feet
length= ? feet

Answer by checkley71(8403)

L=2W+4 IS THE LENGTH
W IS THE WIDTH
W(2W+4)=198
2W^2+4W-198=0
2(W^2+2W-99)=0
2(W+11)(W-9)=0
W-9=0
W=9 FEET IS THE WIDTH.
L=2*9+4=18+4=22 IS THE LENGTH.
PROOF
9*22=198
198=198